OpenStudy (anonymous):

Polar form of complex numbers Which is shown in the graph? 2(cos(-45)+i sin(-45)) or -2(cos(135)+i sin(135))

OpenStudy (anonymous):

I don't get it...don't they equal each other? So how do I know which is the correct one?

OpenStudy (experimentx):

\( r (\cos \theta + i \sin \theta)\) move in the angle \( \theta \), if r is negative ... plot in opposite direction -2(cos(135)+i sin(135)) must be  2(cos(-45)+i sin(-45)) must be  It seems both are equialent